# Geometry of the moduli space of Hermitian-Einstein connections on manifolds with a dilaton

### v1

@@ 5.2 Symmetries of heterotic geometries
$\widehat{R}_{ij}+2\widehat{\nabla}_{i}\partial_{j}\Phi-{1\over 4}P_{ij}=0\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ e^{2\Phi}\widehat{\nabla}_{A}^{i}(e^{-2\Phi}F(C)_{ij})=0\leavevmode\nobreak\ .$ (5.24)
@@
$P_{ij}=-\frac{\alpha^{\prime}}{4}\big{(}\tilde{R}_{ijkm}\tilde{R}_{j}{}^{jkm}-F(C)_{ik\mathrm{a}\mathrm{b}}F_{j}{}^{k\mathrm{a}\mathrm{b}}\big{)}\leavevmode\nobreak\ .$ (5.25)
@@
$dH=\frac{\alpha^{\prime}}{4}\big{(}\tilde{R}^{i}{}_{j}\wedge\tilde{R}^{j}{}_{i}-F(C)^{\mathrm{a}}{}_{\mathrm{b}}\wedge F(C)^{\mathrm{b}}{}_{\mathrm{a}}\big{)}\leavevmode\nobreak\ ,$ (5.26)
with $\tilde{R}$ the curvature of of a connection $\tilde{\nabla}$ of the tangent bundle of $M^{2n}$. Heterotic geometries are more restrictive than those with general $H$ and $P$. In perturbation theory with parameter $\alpha^{\prime}$, $\tilde{R}$ is identified with the curvature $\breve{R}$ of the connection $\breve{\nabla}$ that has torsion $-H$. It is expected that both (5.24) and (5.26) receive higher order corrections in $\alpha^{\prime}$. The result stated above is valid up to the order in $\alpha^{\prime}$ explicitly indicated.
The equations (5.24) with $P$ and $dH$ given as above are the conditions for conformal invariance for the heterotic sigma model or equivalently gradient flow solitons. While, the equations for scale invariance or equivalently for steady flow solitons are
$\widehat{R}_{ij}-\frac{1}{4}P_{ij}=2\widehat{\nabla}_{i}X_{j}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \widehat{\nabla}_{A}^{i}F(C)_{ij}+X^{i}F(C)_{ij}=0\leavevmode\nobreak\ ,$ (5.27)
@@
^$\widehat{\nabla}K=0\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \iota_{K}F(C)=0\leavevmode\nobreak\ ,$ (5.28)
@@
The integrability condition of ^$\widehat{\nabla}K=0$ implies that $\widehat{R}_{ij}{}^{k}{}_{m}K^{m}=0$. Working in the context of perturbation theory and so setting $\tilde{R}=\breve{R}$, see e.g. [46] and reference therein for more explanation in a similar context, at the order considered in $\alpha^{\prime}$, $\breve{R}_{ijkm}=\widehat{R}_{kmij}$, and so $\iota_{K}\breve{R}=0$. This together with (5.26) and the first and second equations in (5.28) imply that $\mathcal{L}_{K}H=0$. Therefore $K$ is $\widehat{\nabla}$-covariantly constant, and so Killing, and leaves $H$ invariant. Thus if $M^{d}$ is compact, it follows from the remark below proposition 2.1 that $K\Phi=0$. In turn such geometries satisfy the condition $D^{i}\big{(}e^{-2\Phi}K_{i}\big{)}=0$ that arises in the context of moduli spaces in the presence of a dilaton field.
@@
$\widehat{R}_{ij}-\frac{1}{4}P_{ij}=2\widehat{\nabla}_{i}\theta_{j}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \widehat{\nabla}_{A}^{j}F(C)_{ji}+\theta^{j}F(C)_{ji}=0\leavevmode\nobreak\ ,$ (5.29)
@@
^$\widehat{\nabla}K=0\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \iota_{K}F(C)=0\leavevmode\nobreak\ ,$ (5.30)
with $K^{\flat}=\theta+2d\Phi$. It is also clear that $\iota_{Y}F(C)=0$ for $Y=-IK$ as $F(C)$ is a $(1,1)$-form. Adapting the results of corollary 5.1 to this case, we can demonstrate the following proposition.
@@
Suppose that $(M^{2n},g,\omega)$ is a CYT manifold with $dH$ given in (5.26), where $F(C)$ is a Hermitian-Einstein connection with $\lambda=0$. Assuming that $(M^{2n},g,\omega)$ is either compact or simply satisfies (5.24), $M^{2n}$ admits two commuting, holomorphic and $\widehat{\nabla}$-covariantly constant vector fields $K$ and $Y=-IK$ with $K^{\flat}=\theta+2d\Phi$ provided that $K^{\flat}\not=0$.
@@
The statement of proposition 5.3 can be adapted to HKT manifolds $(M^{4k},g,\omega_{r})$, as they are special cases of CYT manifolds, provided that $F(C)$ is a $(1,1)$-form with respect to all three complex structures $I_{r}$. The difference is that $K$ is tri-holomorphic and $\widehat{\nabla}$-covariantly constant while $Y_{r}=-I_{r}K$ are holomorphic with respect to $I_{r}$ and $\widehat{\nabla}$-covariantly constant.
@@ 5.3 Application to Hermitian-Einstein connection moduli spaces
Let $(M^{2n},g,\omega)$ be a compact KT manifold that satisfies (5.2) and $K$, with $K^{\flat}=\theta+2d\Phi\not=0$, is a Killing vector field. Then, $\mathscr{M}_{\mathrm{HE}}^{*}(M^{2n})$ admits a strong KT structure and two holomorphic and Killing vector field $\alpha_{K}$ and $\alpha_{Y}$ for $Y=-IK$. In addition, if $(M^{2n},g,\omega)$ is a CYT manifold, then $\mathscr{M}_{\mathrm{HE}}^{*}(M^{2n})$ admits two commuting holomorphic and $\widehat{\mathcal{D}}$-covariantly constant vector field $\alpha_{K}$ and $\alpha_{Y}$.
@@
From the hypothesis of the theorem $K$ is Killing. In addition, proposition 5.1 implies that $K$ is also holomorphic, and $Y=-IK$ is both holomorphic and Killing. As a result the associated vector fields on $\alpha_{K}$ and $\alpha_{Y}$ on the moduli space $\mathscr{M}_{\mathrm{HE}}^{*}(M^{2n})$ are holomorphic and Killing – this follows from theorem 1.2.
If $(M^{2n},g,\omega)$ is a compact CYT manifold that satisfies (5.2), then $K$ and $Y$ are holomorphic and $\widehat{\nabla}$-covariantly constant. As it has already been demonstrated that $\alpha_{K}$ and $\alpha_{Y}$ are holomorphic and Killing because CYT manifolds are a special case of KT manifolds, it remains to demonstrate that $\alpha_{K}$ and $\alpha_{Y}$ are $\widehat{\mathcal{D}}$-covariantly constant. This follows from theorem 1.3 as both $K^{\flat}\wedge K^{\flat}=0$ and $Y^{\flat}\wedge K^{\flat}$ are clearly $(1,1)$-forms. Moreover, $\alpha_{K}$ and $\alpha_{Y}$ commute because $K$ and $Y$ commute. The latter follows because both are $\widehat{\nabla}$-covariantly constant and $K$ is holomorphic, i.e. $i_{K}H$ is a $(1,1)$-form. As a result their commutator is given by $\iota_{K}\iota_{Y}H$ that vanishes. ∎
@@
The Hermitian-Einstein condition (2.7) with $\lambda=0$ on connections of bundles over 4-dimensional manifold implies that the connections are anti-self-dual with respect to the orientation on $M^{4}$ given by the complex structure $I$, i.e. the volume 4-form is chosen as $1/2\,\omega\wedge\omega$, where $\omega$ is the Hermitian form. In fact, the anti-self-duality condition is independent from the choice of complex structure on $M^{4}$. This is the case provided that any two choices of a complex structure induce the same orientation on $M^{4}$. The main consequence of this is that the moduli space of anti-self-dual connections can exhibit a more intricate geometric structure induced by judicious choices of a complex structure on $M^{4}$. Such structures have been explored in [24, 26, 13, 27]. Moreover, there is an additional coincidence in 4-dimensions as the closure of the 3-form torsion $H$ of a KT geometry on $M^{4}$ implies the Gauduchon condition and vice versa. Furthermore, the CYT and HKT structures coincide in four dimensions. In such a case for $M^{4}$ compact and for $H\not=0$, a consequence of the second remark below corollary 5.2 implies that $M^{4}$ must be identified with $S^{3}\times S^{1}$ up to an identification with a discrete group. The properties of the moduli spaces of anti-self-dual connections on $S^{3}\times S^{1}$ have been extensively been explored in [13], see also [27].
It remains to consider the geometry of instanton moduli spaces for $(M^{4},g,\omega)$ a compact KT manifold. This has already been dealt with in theorem 5.1. It should also be noted that if $M^{4}$ admits an oriented bi-KT structure, or equivalently oriented generalised Kähler [24, 25], then the analogue of the condition (5.1) associated to second KT structure with connection $\breve{\nabla}$ that has torsion $-H$ is
$D^{i}\big{(}e^{-2\Phi}(-\theta_{i}+2\partial_{i}\Phi)\big{)}=0\leavevmode\nobreak\ .$ (5.31)

### v2

@@ 5.2 Symmetries of heterotic geometries
$\widehat{R}_{ij}+2\widehat{\nabla}_{i}\partial_{j}\Phi-{1\over 4}P_{ij}=0\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ e^{2\Phi}\widehat{\nabla}_{A}^{i}(e^{-2\Phi}F(C)_{ij})=0\penalty 10000\ .$ (5.24)
@@
[…]
@@
[…]
The equations (5.24) with $P$ and $dH$ given as above are the conditions for conformal invariance for the heterotic sigma model or equivalently gradient flow (generalised) solitons. While, the equations for scale invariance or equivalently for steady flow (generalised) solitons are
$\widehat{R}_{ij}-\frac{1}{4}P_{ij}=2\widehat{\nabla}_{i}X_{j}\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \widehat{\nabla}_{A}^{i}F(C)_{ij}+X^{i}F(C)_{ij}=0\penalty 10000\ ,$ (5.27)
@@
$\widehat{\nabla}K=0\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ \iota_{K}F(C)=0\penalty 10000\ ,$ (5.28)
@@
The integrability condition of $\widehat{\nabla}K=0$ implies that $\widehat{R}_{ij}{}^{k}{}_{m}K^{m}=0$. Working in the context of perturbation theory and so setting $\tilde{R}=\breve{R}$, see e.g. [47] and reference therein for more explanation in a similar context, at the order considered in $\alpha^{\prime}$, $\breve{R}_{ijkm}=\widehat{R}_{kmij}$, and so $\iota_{K}\breve{R}=0$. This together with (5.26) and the first and second equations in (5.28) imply that $\mathcal{L}_{K}H=0$. Therefore $K$ is $^\widehat{\nabla}$-covariantly constant, and so Killing, and leaves $H$ invariant. Thus if $M^{d}$ is compact, it follows from the remark below Proposition 2.1 that $K\Phi=0$. In turn such geometries satisfy the condition $D^{i}\big(e^{-2\Phi}K_{i}\big)=0$ that arises in the context of moduli spaces in the presence of a dilaton field.
@@
$\widehat{R}_{ij}-\frac{1}{4}P_{ij}=2\widehat{\nabla}_{i}\theta_{j}\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \widehat{\nabla}_{A}^{j}F(C)_{ji}+\theta^{j}F(C)_{ji}=0\penalty 10000\ ,$ (5.29)
@@
$\widehat{\nabla}K=0\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ \iota_{K}F(C)=0\penalty 10000\ ,$ (5.30)
with $K^{\flat}=\theta+2d\Phi$. It is also clear that $\iota_{Y}F(C)=0$ for Y=IKY=-IK as $F(C)$ is a $(1,1)$-form. Adapting the results of Corollary 5.1 to this case, we can demonstrate the following proposition.
@@
Suppose that $(M^{2n},g,\omega)$ is a CYT manifold with dHdH given in (5.26), where $F(C)$ is a Hermitian-Einstein connection with $\lambda=0$. Assuming that $(M^{2n},g,\omega)$ is either compact or simply satisfies (5.24), $M^{2n}$ admits two commuting, holomorphic and $^\widehat{\nabla}$-covariantly constant vector fields $K$ and Y=IKY=-IK with $K^{\flat}=\theta+2d\Phi$ provided that $K^{\flat}\not=0$.
@@
The statement of Proposition 5.3 can be adapted to HKT manifolds $(M^{4k},g,\omega_{r})$, as they are special cases of CYT manifolds, provided that $F(C)$ is a $(1,1)$-form with respect to all three complex structures $I_{r}$. The difference is that $K$ is tri-holomorphic and $^\widehat{\nabla}$-covariantly constant while $Y_{r}=-I_{r}K$ are holomorphic with respect to $I_{r}$ and $^\widehat{\nabla}$-covariantly constant.
@@ 5.3 Application to Hermitian-Einstein connection moduli spaces
Let $(M^{2n},g,\omega)$ be a compact KT manifold that satisfies (5.2) and $K$, with $K^{\flat}=\theta+2d\Phi\not=0$, is a Killing vector field. Then, $\mathscr{M}_{\mathrm{HE}}^{*}(M^{2n})$ admits a strong KT structure and two holomorphic and Killing vector field $\alpha_{K}$ and $\alpha_{Y}$ for Y=IKY=-IK. In addition, if $(M^{2n},g,\omega)$ is a CYT manifold, then $\mathscr{M}_{\mathrm{HE}}^{*}(M^{2n})$ admits two commuting holomorphic and $\widehat{\mathcal{D}}-$covariantly constant vector field $\alpha_{K}$ and $\alpha_{Y}$.
@@
From the hypothesis of the theorem $K$ is Killing. In addition, Proposition 5.1 implies that $K$ is also holomorphic, and Y=IKY=-IK is both holomorphic and Killing. As a result the associated vector fields on $\alpha_{K}$ and $\alpha_{Y}$ on the moduli space $\mathscr{M}_{\mathrm{HE}}^{*}(M^{2n})$ are holomorphic and Killing – this follows from Theorem 1.2.
If $(M^{2n},g,\omega)$ is a compact CYT manifold that satisfies (5.2), then $K$ and $Y$ are holomorphic and $^\widehat{\nabla}$-covariantly constant. As it has already been demonstrated that $\alpha_{K}$ and $\alpha_{Y}$ are holomorphic and Killing because CYT manifolds are a special case of KT manifolds, it remains to demonstrate that $\alpha_{K}$ and $\alpha_{Y}$ are $\widehat{\mathcal{D}}-$covariantly constant. This follows from Theorem 1.3 as both $K^{\flat}\wedge K^{\flat}=0$ and $Y^{\flat}\wedge K^{\flat}$ are clearly $(1,1)$-forms. Moreover, $\alpha_{K}$ and $\alpha_{Y}$ commute because $K$ and $Y$ commute. The latter follows because both are $^\widehat{\nabla}$-covariantly constant and $K$ is holomorphic, i.e. $i_{K}H$ is a $(1,1)$-form. As a result their commutator, given by $\iota_{K}\iota_{Y}H$, vanishes. ∎
@@
The Hermitian-Einstein condition (2.7) with $\lambda=0$ on connections of bundles over 4-dimensional manifold implies that the connections are anti-self-dual with respect to the orientation on $M^{4}$ given by the complex structure $I$, i.e. the volume 4-form is chosen as $1/2\,\omega\wedge\omega$, where $\omega$ is the Hermitian form. In fact, the anti-self-duality condition is independent from the choice of complex structure on $M^{4}$. This is the case provided that any two choices of a complex structure induce the same orientation on $M^{4}$. The main consequence of this is that the moduli space of anti-self-dual connections can exhibit a more intricate geometric structure induced by judicious choices of a complex structure on $M^{4}$. Such structures have been explored in [24, 26, 13, 27]. Moreover, there is an additional coincidence in 4-dimensions because the closure of the 3-form torsion $H$ of a KT geometry on $M^{4}$ implies the Gauduchon condition and vice versa. Furthermore, the CYT and HKT structures coincide in four dimensions. In such a case for $M^{4}$ compact and for $H\not=0$, a consequence of the second remark below Corollary 5.2 implies that $M^{4}$ must be identified with $S^{3}\times S^{1}$ up to an identification with a discrete group. The properties of the moduli spaces of anti-self-dual connections on $S^{3}\times S^{1}$ have been extensively been explored in [13], see also [27].
[…]
$D^{i}\big(e^{-2\Phi}(-\theta_{i}+2\partial_{i}\Phi)\big)=0\penalty 10000\ .$ (5.31)